# MoVE-KD: Knowledge Distillation for VLMs with Mixture of Visual Encoders

### v1

@@ Abstract
Visual encoders are fundamental components in vision-language models (VLMs), each showcasing unique strengths derived from various pre-trained visual foundation models. To leverage the various capabilities of these encoders, recent studies incorporate multiple encoders within a single VLM, leading to a considerable increase in computational cost. In this paper, we present Mixture-of-Visual-Encoder Knowledge Distillation (MoVE-KD), a novel framework that distills the unique proficiencies of multiple vision encoders into a single, efficient encoder model. Specifically, to mitigate conflicts and retain the unique characteristics of each teacher encoder, we employ low-rank adaptation (LoRA) and mixture-of-experts (MoEs) to selectively activate specialized knowledge based on input features, enhancing both adaptability and efficiency. To regularize the KD process and enhance performance, we propose an attention-based distillation strategy that adaptively weighs the different visual encoders and emphasizes valuable visual tokens, reducing the burden of replicating comprehensive but distinct features from multiple teachers. Comprehensive experiments on popular VLMs, such as LLaVA and LLaVA-NeXT, validate the effectiveness of our method. The code will be released.
@@ 1 Introduction
Figure 1: Comparison of LLaVA-1.5-7B [25] and RADIO [34] on a wide range of 8 benchmarks, and MoVE-KD surpasses them on all benchmarks.
@@ 1.7B Models
+ MoVE-KD (Ours) MobileLLaMA-1.4B [8] 72.9 56.6 43.4 32.1 84.8 56.1 1182.3 47.4
@@
+ MoVE-KD (Ours) Vicuna-7B [51] 79.5 63.2 58.3 52.3 86.9 69.3 1524.5 66.3
@@
+ MoVE-KD (Ours) Vicuna-7B [51] 82.3 64.5 63.7 58.0 86.7 70.7 1537.2 67.6
@@
+ MoVE-KD (Ours) Vicuna-13B [51] 80.6 64.2 59.7 55.7 85.7 73.2 1568.1 70.2
@@
+ MoVE-KD (Ours) Vicuna-13B [51] 83.1 65.7 65.8 60.9 86.8 73.7 1579.3 70.6
@@ 3.3 Attention-guided KD regularization
Teacher weight. For the coarse teacher-level regularization, we take the Softmax of the mean value of the cross-attention between the [CLS] token ${V^{(cls)}}$ and the $i$-th teacher’s tokens ${V^{(t)}_{i}}\in\mathbb{R}^{n\times d}$ as its weight. These weights of various teachers, indicate the responses of teachers to the specific images, and thus showing their contributions to the recognition. Mathematically, the teacher-level regularization is computed as:
@@ 4.3 Ablation study
We present the results of the ablation study for each design in our method in Tab. 2. For simplicity, we use the LLaVA1.5-7b model as the baseline and incrementally add each method design from MoVE-KD.
@@
In addition, since the student encoder is initialized by the pre-trained CLIP, as shown in Tab. 4, we find that it is also necessary to give regularization with a higher weight to the CLIP teacher encoder in practice. At the same time, due to the differences in the capabilities of different encoders, if the weights of the teachers are simply uniformized, it cannot allow each teacher to fully exert its strengths in its area of expertise. The teacher weight flexibly allocates the weights of each teacher according to the input, which indeed can improve the performance to a certain extent. Therefore, the thinking about the optimal method of teacher weight allocation will continue. In Fig. 5, we show the weight distribution of each teacher across different images.

### v2

@@ Abstract
Visual encoders are fundamental components in vision-language models (VLMs), each showcasing unique strengths derived from various pre-trained visual foundation models. To leverage the various capabilities of these encoders, recent studies incorporate multiple encoders within a single VLM, leading to a considerable increase in computational cost. In this paper, we present Mixture-of-Visual-Encoder Knowledge Distillation (MoVE-KD), a novel framework that distills the unique proficiencies of multiple vision encoders into a single, efficient encoder model. Specifically, to mitigate conflicts and retain the unique characteristics of each teacher encoder, we employ low-rank adaptation (LoRA) and mixture-of-experts (MoEs) to selectively activate specialized knowledge based on input features, enhancing both adaptability and efficiency. To regularize the KD process and enhance performance, we propose an attention-based distillation strategy that adaptively weighs the different encoders and emphasizes valuable visual tokens, reducing the burden of replicating comprehensive but distinct features from multiple teachers. Comprehensive experiments on popular VLMs, such as LLaVA and LLaVA-NeXT, validate the effectiveness of our method. Our code is available at: https://github.com/hey-cjj/MoVE-KD.
@@ 1 Introduction
Figure 1: Comparison of LLaVA-1.5-7B [26] and RADIO [35] on a wide range of benchmarks, and MoVE-KD surpasses them.
@@ 1.7B Models
+ MoVE-KD-v1.0 MobileLLaMA-1.4B [8] 72.9 56.6 43.4 32.1 84.8 56.1 1182.3 47.4
+ MoVE-KD-v1.1 MobileLLaMA-1.4B [8] 73.8 57.7 44.3 29.3 86.1 57.3 1188.4 48.8
@@
+ MoVE-KD-v1.0 Vicuna-7B [52] 79.5 63.2 58.3 52.3 86.9 69.3 1524.5 66.3
+ MoVE-KD-v1.1 Vicuna-7B [52] 79.9 63.9 59.6 52.7 86.3 69.8 1509.1 67.4
@@
+ MoVE-KD-v1.0 (Ours) Vicuna-7B [52] 82.3 64.5 63.7 58.0 86.7 70.7 1537.2 67.6
@@
+ MoVE-KD-v1.0 Vicuna-13B [52] 80.6 64.2 59.7 55.7 85.7 73.2 1568.1 70.2
+ MoVE-KD-v1.1 Vicuna-13B [52] 80.8 63.9 61.1 57.5 86.3 71.8 1568.3 69.7
@@
+ MoVE-KD-v1.0 (Ours) Vicuna-13B [52] 83.1 65.7 65.8 60.9 86.8 73.7 1579.3 70.6
@@ 3.3 Attention-guided KD regularization
Teacher weight. For the coarse teacher-level regularization, we take the Softmax of the mean value of the cross-attention between the [CLS] token ${V^{(cls)}}$ and the $i$-th teacher’s tokens ${V^{(t)}_{i}}\in\mathbb{R}^{n\times d}$ as its weight. These weights of various teachers, indicate the responses of teachers to the specific images, and thus showing their contributions to the recognition, which is formulated as:
@@ 4.3 Ablation study
We present the results of the ablation study for each design in our method in Tab. 2. For simplicity, we use the LLaVA-1.5 7b model as the baseline and incrementally add each method design from MoVE-KD.
@@
Since the student encoder is initialized by the pre-trained CLIP, as shown in Tab. 4, we find that it is necessary to give a higher weight to the CLIP teacher encoder in practice. What’s more, due to the differences in the capabilities of different encoders, if the weights of the teachers are simply uniformized, it cannot allow each teacher to fully exert its strengths in its area of expertise. The teacher weight flexibly allocates the weights of each teacher according to the input, which indeed can improve the performance to a certain extent. However, the optimal method of teacher weight allocation needs further exploration.
@@ 6 Conclusion
Acknowledgement
This work was supported by the National Natural Science Foundation of China (62476011).